Home‎ > ‎Reports‎ > ‎Crew103‎ > ‎

day06


Crew 103 Log Book for April 14, 2011

Photos of the day





Commander's Check-In Report

Anne-Sophie Kalbfleisch Reporting

Crew Physical Status: everybody is fine. just still have some headache after the accident

Time Departed/Returned from EVA:

First EVA 11h30-12h30

Second EVA 15h30-16h30

Brief Narrative of Field Mission Results:

We tried to get more comfortable with the max rover. We were divided in two teams, one that stayed in the station and the other that was outside. We tried to conduct it better and learned how to communicate with each other in order to facilitate the rover's conduction. During the afternoon we began our experience with the algae. We putted some droplets on an agar-filled "Petri" box, and putted them on our LEDs electronic device. Each day we will take some pictures in order to check their development.

EVA Data/Interpretations:

Because the battery capacity of the max rover is limited, we learned during the first EVA that it could be better to take some extra battery with us during the EVA. That's what we did during the second EVA. For unknown reasons the rover didn't work anymore, even after replacing his battery. We will check this out tomorrow.

Engineering/Hab Maintenance:

Nothing to mention

Plans for Tomorrow:

We will probably begin a new geological experience. In continuity of the previous crew, we will try to determine the permeability of the surrounding soils. Because of the high variety of different soils, they just analyzed two different kind of soils. We will make a more exhaustive analysis by collecting and analyzing much more soils.

Inventory:

Nothing to mention

Miscellaneous:

Don Lusko sent a fax with some informations to Fed-ex. Let's hope it will work



Commander's Report

Anne-Sophie Kalbfleisch Reporting

Today we made two EVA's. The aim was to train different skills related with the max rover : conducting it at distance, communicating important information through the walkie talkie, checking what's missing or failing for a long term mission, changing a battery and so on. The second EVA was a second trial in order to improve this interaction.

During the late afternoon we started our experience with the algae. In the Science Report you will get more information about the protocol followed and what we did precisely. We also spent some time replying to the forms sent by the psychologist who is following us.


Engineering Report

Thibault Windels Reporting

Generator/Electricity:

Honey (Diesel Generator) (Used? Yes/No): yes

Wendy (Gas Generator) (Used? Yes/No): no

Battery Charge Level: 10,3 volts

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0,5

Propane Tank (Full/0.75/0.5/0.25/Empty): 0,40

Gasoline Tank (Full/0.75/0.5/0.25/Empty): The same as yesterday, we did not use gasoline today

Notes/Comments: /


Oil:

Kitty Oil Quantity (# of Quarts in storage): the previous commander told us not to check the oil quantity of Kitty because we have to turn the all generator off!

Honey Oil Quantity (# of Quarts in storage): /

ATV Oil Quantity (# of Quarts in storage): full

Car Oil Quantity (# of Quarts in storage): /


Water Status:

Outside Potable Water Tank Level (inches from bottom): 14,5

Trailer Potable Water Tank Level (inches from bottom): 14,5

Hab Potable Water Tank Level (inches from bottom): 10,5

Potable Water Meter Reading: 45123,6

In to GreenHab Meter Reading: /

Notes/Comments: /


GreenHab:

Condition of plants in Tank 1: /

Condition of plants in Tank 2: /

Condition of plants in Tank 3: /

Greenhab Notes/Comments: The ventilation of the green hab is off. The thermostat is at 50°F and the temperature inside is 77°F


Observatory:

Telescope Used? (Yes/No): no

Observatory Notes/Comments: there's no telescope


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): yes

Oil Checked (yes/no): yes

Tire Status: ok

Odometer at end of day: 124234,5

Notes/Comments on Hab Car: We used the car because we had to call Fed Ex with Don Lusko.


ATV 1 (Opportunity):

Used: No

Oil Checked: no

Tire Status: ok

Comments: /


ATV 2 (Spirit):

Used: no

Oil Checked: no

Fuel Consumed: 0

Tire Status: ok

Comments: /


ATV 3 (Viking I):

Used: no

Oil Checked: no

Fuel Consumed: 0

Tire Status: ok

Comments: /


ATV 4 (Viking II):

Used: no

Oil Checked: no

Fuel Consumed: 0

Tire Status:

Comments: does not work


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 65

Actual temperature upstairs (Farenheit): 71

Thermostat setting downstairs (Farenheit): /

Actual temperature downstairs (Farenheit): 64


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: internet is still running well


General Engineering Notes/Comments:



Science Report (Biology)

Wouter de Groot Reporting

Today we started the LED experiment where the algae are growing on a solid substrate.

As I told you yesterday (without writing a science report), we have prepared a basic solution containing a low algae concentration. Today we have displayed sixteen small quantities of this basic solution (20 micro-liters, for that we use the biological material that is available in the  Hab lab) at regular space intervals on an agare substrate that we have placed on a "home made" circuit containing different types (colors) of LEDs. It is difficult to explain exactly how is this circuit but I will send you a picture of it tomorrow so you will understand easier what I am talking about. These LEDs will be working all the time during the experiment.

Since this experiment is more qualitative than quantitative, we will simply take photographs at regular intervals, one around midnight (the first one), the second around 8am, then around 4 pm and so on, every 8 hours. Depending on which LED's color the algae are, we will see which ones will grow the quickest.

I don't know exactly what do you expect as science report. If you want more details, ask me!


EVA 1 Report

Anne-Sophie Kalbfleisch Reporting

Time of Departure:  11H30

Time of Arrival: 12H30

Crew Members: J. Offergeld, G. Van Roy

Site Location: Just in front of the Hab

Transit Mode: By foot

Objectives: The purpose of this EVA was to increase our interaction with the max rover as well as with the crew members stayed inside the Hab.

Results:  A good training with the joystick is really important because otherwise the rover doesn't obey appropriately.

Lessons Learned: It's better to take extra batteries with us.


EVA 2 Report

Anne-Sophie Kalbfleisch Reporting

Time of Departure: 15H30

Time of Arrival: 16H30

Crew Members: T. Windels, A. Kalbfleisch, W. de Groot

Site Location: just in front of the Hab

Transit Mode: By foot

Objectives: The purpose of this EVA was to increase our interaction with the max rover as well as with the crew members stayed inside the Hab. This time we took an extra battery with us and fixed a camera above the rover.

Results:  the camera worked properly, but the rover stopped after a short period and didn't start again even after having replaced the battery.

Lessons Learned: we will try to find out what failed.


Chef's Report

Gregory Van Roy Reporting

1a) Today was

A cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

Nothing special today

2)  List the foods served at today's main meal (usually the evening meal), giving full names of each.

Cup a Soup Chicken, Beef and Pepperoni Pizza, Bread, Kool-Aid drink mix

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

Cup a Soup : left 1 pack

Beef : left 1 pack

4)  Recipes for anything more complicated than following package directions exactly, or rehydrating.

Pizza : 3 ¾ cup of flour, ¾ TBSP salt, 1 ½ TBSP oil, 1 ½ cup of water

+ pepperoni, tomato sauce, spices

+ beef, cheddar, tomato sauce, spices

Brioche : 9 TBSP butter, 3 cup of flour, 3 TBSP sugar, ¾ TBSP salt, 2

eggs, 1 ½ dry yeast, ¾ cup of water

5)  Lessons learned  (if any)

We have a weird taste for the brioche, we will try to put more sugar next time.

6) Comments/questions for Kim and Jean, the food study investigators

We consume a lot of milk the morning with the Qwaker and Corn Flakes as well as Coffee and Hot Chocolate… For the hot chocolate, it’s possible to use water instead of milk but the taste is really better with milk for Qwaker! We already consumed more than a half of the reserve we have at the beginning!

We discover a huge stock of tea and tofu… I think the previous crews doesn’t need so much quantity… It’s pretty the same for the sugar.

Otherwise, we use a lot of flour as well with cooking bread, pancakes, brioche, pizza,…